[1,1] [0,3]
2x+y=3

Answer by fanism(39)

the equation is:
x2-x1/y2-y1 = x-x1/y-y1
in your case,
x1, y1 = [1,1]
x2, y2 = [0,3]
therefore:
0-1/3-1 = x-1/y-1
-1/2 = x-1/y-1
-y+1=2x-2
0=2x+y-3